ROBOTS CAN BE MADE WITH HUMAN-LIKE EMOTIONS—BUT SHOULDN’T BE John McCarthy Computer Science Department Stanford University jmc@cs.stanford.edu http://www-formal.stanford.edu/jmc/ March 22, 2004

A slogan for AI: Whatever a person can do, the person should be able to make a computer do for him.

Almost all of my papers are on the web page.

USEFUL ROBOTS WON’T HAVE HUMAN-LIKE EMOTIONS

• human motivations and emotion —–direct reaction to inputs —–chemical state —–subgoals escaping from main goal

• Robot and Baby story http://www.formal.stanford.edu/jmc/robotandbaby.html • hatred. Police robots shouldn’t hate criminals.

• depression, joy - I can’t do them now, but maybe you can.

• robots in society. Robots shouldn’t be capable of suffering or personal ambition.

• Computer understanding emotion requires logical formalization.

ONE PERSON HATES ANOTHER

• If person1 hates person2, person1 prefers situations in which person2 is worse off.

• Hates(person1, person2, s)

• Hates(person1, person2, s)∧W orseOff(person2, Result(e, s), s) → P refers(person1, Result(e, s), s)